OpenStudy (anonymous):

Help please What are the zeroes of the function? Graph the function. y = x(x+2)(x+5) a. 0,2,5 b,-2,-5 c -2,-5, 2 d 0,-2, -5

OpenStudy (anonymous):

zeroes of the function is where y becomes 0. now we know that anything multiplied by 0 is zero right?

OpenStudy (anonymous):

so to make y = 0 in our equation, we can have either x = 0, or (x+2) = 0 or (x+5) = 0 right?

OpenStudy (anonymous):

so would it be a? Sorry for late response @SandeepReddy

OpenStudy (anonymous):

no if x= 0, then obv. x = 0 but if x+2 = 0 then x +2 -2 = 0 -2 => x = -2 ( just to be shortcut, plus becomes minus if we jump value from either side of equals , here +2 became -2 ) so x = -2 similarly x+5 = 0 -> x = -5 so x = 0, -2, -5

OpenStudy (anonymous):

ohh okay can you help with one mor

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what are the zeros of the function? what are their multiplicities? f(x)=x^4-4x^3+3x^2

OpenStudy (anonymous):

A. The numbers -1 and -3 are zeroes of multiplicities2;the numbers 1 and 3 are zeroes of multiplicities 1. B. the number 0 is a zero of multiplicity2;the numbers 1 and 3 are zeroes of multiplicty 1. C. The numbers 0 and 1 are the zeroes of multpilcty2; the number 3 is a zero of multiplcty1. D.The number 0 is zero of multiplicty2;the numbers -1 and -3 are zeroes of multiplicty

OpenStudy (anonymous):